# Repeated Padding for Sequential Recommendation

### v2

@@ 2.4. Ours: Repeated Padding
Table 2. Comparison of data augmentation between our RepPad and several representative methods.
Methods Augmentation Type No Hyper-Parameters No Training Model Agnostic Total Training Sequences
Only Augmentation (Tan et al., 2016a; Tang and Wang, 2018) Random or Elaborate Operators $\geq$ 2$\left|\mathcal{U}\right|$
ASReP (Liu et al., 2021b), BiCAT (Jiang et al., 2021) Bidirectional Transformers $>$ $\left|\mathcal{U}\right|$
RepPad (Ours) Repeated Padding $=$ $\left|\mathcal{U}\right|$
In order to achieve this objective, we propose a simple but effective padding strategy called Repeated Padding (RepPad). Specifically, we utilize the original sequence as padding content instead of the special value 0. As shown in Figure 1 (d), give the original sequence $s_{u}=[v_{1},v_{2},v_{3},v_{4}]$, and we pad the original sequence once on the left and get the padded sequence $s_{u}^{p}=[\underline{v_{1},v_{2},v_{3},v_{4}},v_{1},v_{2},v_{3},v_{4}]$. This operation can be formulated as:
where $m$ is padding times and |||| indicates sequence concatenation. This operation can be performed only once or repeated several times until there is not enough space left to hold the original sequence. In other words, $m$ can be pre-specified or calculated based on the original sequence length $s_{u}$ and the maximum sequence length $N$. For example, when the maximum sequence length is 10, the remaining length of 2 is insufficient to accommodate the original sequence of length 4, so it still needs to apply function Eq. (2). The final input sequence $s_{u}^{f}=[0,0,\underline{v_{1},v_{2},v_{3},v_{4}},v_{1},v_{2},v_{3},v_{4}]$. In addition, given the original sequence $s_{u}=[v_{3},v_{7},v_{9}]$, the padding operation can be performed up to 3 times, and the final sequence $s_{u}^{f}=[0,0,\underline{v_{3},v_{7},v_{9}},\underline{v_{3},v_{7},v_{9}},v_{3},v_{7},v_{9}]$.
However, there is a problem with the above padding strategy: there may be cases where the end of the original sequence is used to predict the head. Using $[0,0,\underline{v_{1},v_{2},v_{3},v_{4}},\underline{v_{1}},v_{2},v_{3},v_{4}]$ as training data, there will be cases where $v_{1},v_{2},v_{3},v_{4}$ is used to predict $v_{1}$. To solve this problem, we add the special value 0 between each repeated padding to prevent this from happening. This operation can be formulated as:
[…]
As shown in Figure 1 (e), given original sequence $s_{u}=[v_{1},v_{2},v_{3},v_{4}]$, the final sequence after Eq. (4) and (2) is $[0,v_{1},v_{2},v_{3},v_{4},0,v_{1},v_{2},v_{3},v_{4}]$. We give the possible values of $m$ and their interpretation in Table 1.
Our method has no trainable parameters or hyper-parameters that need to be manually tuned. It is a model-agnostic plug-and-play data augmentation approach. In the Experiments section, we will explore and analyze several variants of RepPad, including different times of padding and whether to add the delimiter 0. The best-performing variant is The best-performing variant is randomly selected in the range from one to the maximum number of times. Our method is very easy to implement, and we present its Python procedure in Algorithm 1. So far, we have introduced the basic idea of our RepPad. It is worth emphasizing that the repeated padding sequences still need to apply function Eq. (2) operation before they can be fed into the model:
(5) $s_{u}^{f}=SeqPrepare\left(s_{u}^{p}\right).$
2.5. Discussions
2.5.1. Comparison with Existing Methods.
The advantages of our approach are as follows: (i) No increase in number of sequences: From Figure 1, we can observe that our approach performs data augmentation without increasing the amount of data for training. While previous approaches can be considered as extending the length of the dataset vertically, our approach utilizes the free space horizontally. For traditional data augmentation methods, the amount of data involved in training will be at least 2$\left|\mathcal{U}\right|$. For sequential recommendation models that leverage both data augmentation and contrastive learning, each sequence is augmented at least twice, and with the addition of the original sequence, the number of model sequences involved in training is at least 3$\left|\mathcal{U}\right|$. (ii) No Hyper-parameters: These existing augmentation methods or augmentation operators (Slide Window, Dropout, Crop, Insert, Mask, Substitute, Reorder, and so on.) typically contain at least one hyper-parameter for each that must be manually tuned based on the dataset. However, our approach does not involve any hyperparameters that need to be tuned carefully. (iii) No Training Cost and Model Agnostic: Some researchers propose to use bidirectional Transformers (Liu et al., 2021b; Jiang et al., 2021) or diffusion models (Liu et al., 2023) for data augmentation. These methods require training of data augmentation modules in addition to training of backbone recommendation networks and have limitations on the type of backbone network. In contrast, our method does not contain any parameters, does not require a training process, and can be adapted to most sequential models. We present a summary comparison between PepPad and existing methods in Table 2. It should be emphasized that our focus is on sequence-level data augmentation, which is also the most mainstream practice in data augmentation in the field of sequential recommendation. Augmentation at the embedding (Qiu et al., 2022) and model (Hao et al., 2023) levels are not involved in our discussion.
@@ 2.5. Discussions
max_pad_num = int(max_len / (len(orig_seq[:-3]) + 1))
# Randomly select a padding count m.
@@
# Perform the Repeated Padding operation to get s_u^p.
for i in range(pad_num):
final_input = orig_seq[:-3] + [0] + final_input
final_target = orig_seq[1:-2] + [0] + final_target
@@
Our approach intends to maximize the use of idle input space. Intuitively, RepPad is more effective on short-sequence datasets and may be ineffective on long-sequence datasets. This is because for long-sequence datasets, most users have sequence lengths that close to or exceed the given maximum sequence length, and thus, there will be no space left for padding. We will verify our conjecture in the experimental section.
@@ 3.1. Experiment Setup
3.2. Datasets
The experiments are conducted on six datasets:
@@ 3.2. Datasets
  • Toys, Beauty, Sports, and Home: Four datasets obtained from Amazon (McAuley et al., 2015) and correspond to the ”Toys and Games”, ”Beauty”, ”Sports and Outdoors”, and ”Home” categories, respectively.
  • Yelp [^1] https://www.yelp.com/dataset: It is the second version of the Yelp dataset released in 2020. We utilize the records after January 1st, 2019.
@@
  • MovieLens-1M [^2] https://grouplens.org/datasets/movielens/ (ML-1M for short): This dataset comprises rating behaviors gathered from a movie review website.
@@
The Amazon and Yelp datasets contain mainly short user sequences. The MovieLens-1M contains mainly long user sequences. Following (Liu et al., 2021a; Dang et al., 2023a; Tian et al., 2023), we reduce the data by extracting the 5-core. The maximum sequence length $N$ is set to 200 for the MovieLens-1M dataset and 50 for other datasets. The detailed statistics are summarized in Table 3.
Table 3. The statistics of six datasets. The Inter donates the number of interactions and AvgLen donates average length.
Dataset Toys Beauty Sports Yelp Home ML-1M
# Users 19,412 22,363 35,958 30,431 66,519 6,040
# Items 11,924 12,101 18,357 20,033 28,237 3,706
# Inter 167,597 198,502 296,337 316,354 551,682 1,000,209
# AvgLen 8.6 8.9 8.3 10.4 8.3 165.6
Sparsity 99.93% 99.92% 99.95% 99.95% 99.97% 95.53%
3.3. Baselines
@@
In order to conduct a comprehensive evaluation of the effectiveness and superiority of RepPad, the baselines consist of three main categories:
@@
First. The first category is eight representative sequential recommendation models, which are used to evaluate whether RepPad is effective in improving the model’s performance.
@@
  • RNN-based models: GRU4Rec (Hidasi et al., 2015b) leverages gated recurrent units to capture behavioral patterns. NARM (Li et al., 2017) incorporates attention into RNN to model the user’s sequential pattern.
@@
  • CNN-based models: Caser (Tang and Wang, 2018) embeds a sequence of recent items in the time and latent spaces, then learns sequential patterns using convolutional filters. NextItNet (Yuan et al., 2019) combines masked filters with 1D dilated convolutions to model the dependencies.
  • MLP-based models: FMLP-Rec (Zhou et al., 2022) is an all-MLP model with learnable filters for sequential recommendation tasks.
  • Transformer-based models: SASRec (Kang and McAuley, 2018) leverages the multi-head attention mechanism to perform sequential recommendation. LightSANs (Fan et al., 2021) proposes the low-rank decomposed self-attention networks to overcome quadratic complexity and vulnerability to over-parameterization. STOSA (Fan et al., 2022) embeds each item as a stochastic Gaussian distribution and devises a Wasserstein Self-Attention module to characterize item-item relationships.
@@
Second. The second category is the widely used stochastic data augmentation methods, which do not require training but include several hyper-parameters that need to be manually tuned.
@@
  • Slide Window (SW) (Tan et al., 2016b): It adopts slide windows to intercept multiple new subsequences from the original sequence.
@@
  • CMR (Xie et al., 2022): This work proposes three sequence data augmentation operators with contrastive learning. In this category, we only use the three operators including Crop, Mask, and Reorder.
@@
  • CMRSI (Liu et al., 2021a): Based on CMR, this work proposes two additional operators with contrastive learning. We adopt five operators including Crop, Mask, Reorder, Substitute, and Insert.
@@
Third. The third category is representative of training-required data augmentation models, which usually include trainable parameters and hyper-parameters that need to be manually tuned.
@@
  • ASReP (Liu et al., 2021b): This method employs a reversely pre-trained transformer to generate pseudo-prior items for short sequences. Then, fine-tune the pre-trained transformer to predict the next item.
@@
  • DiffuASR (Liu et al., 2023): It adopts the diffusion model for sequence generation. Besides, two guide strategies are designed to control the model to generate the items corresponding to the raw data.
@@
  • CL4SRec (Xie et al., 2022): This method leverages random data augmentation and utilizes contrastive learning to extract self-supervised signals from the original data.
@@
We do not include method L2Aug (Wang et al., 2022) and CASR (Wang et al., 2021) since they do not provide open-source codes for reliable reproduction.
3.4. Evaluation Metrics
@@
Following (Liu et al., 2021a; Dang et al., 2023a; Tian et al., 2023), we adopt the leave-one-out strategy, wherein the last item of each user sequence serves as the test data, the items preceding it as validation data, and the remaining data as training data. We rank the prediction over the whole item set rather than negative sampling, otherwise leading to biased discoveries (Krichene and Rendle, 2020). The evaluation metrics include Hit Ratio@K (denoted by HR@K), and Normalized Discounted Cumulative Gain@K (NDCG@K). We report results with K $\in\{5,10,20\}$. Generally, greater values imply better ranking accuracy.
3.5. Implementation Details

### v3

@@ 3.4. The Previous Work: Repeated Padding (RepPad)
max_pad_num = int(max_len / len(final_input))
# Randomly select a padding count.
@@
# Perform the Repeated Padding operation.
final_input = (final_input + [0]) * pad_num + final_input
final_target = (final_target + [0]) * pad_num + final_target
@@
The Python procedure of RepPad is presented in Algorithm 1. So far, the basic idea of RepPad has been introduced. It is worth emphasizing that the repeated padding sequences still need to apply function Eq. (2) operation before they can be fed into the model:
(5) $s_{u}^{f}=SeqPrepare\left(s_{u}^{p}\right).$
3.5. RepPad+: Augmenting the Medium-length Sequences
In Figure 1 (d) and (e), we can observe that sequence $[0,0,0,0,v_{6},v_{4},v_{1},v_{7},v_{9},v_{2}]$ cannot be augmented because of insufficient remaining space. However, for this sequence, there is still idle input space of length 4. Naturally, we can consider padding this space using consecutive sub-sequences of the original sequence. However, if we follow the general interception paradigm, i.e., only the most recent interactions are intercepted, it may cause earlier interactions to be ignored in training, resulting in inaccurate representations learning. Inspired by Sliding Windows (Tang and Wang, 2018), We propose to randomly truncate subsequences from the original sequence for padding. Specifically, we first calculate the length of the remaining space based on the maximum sequence length $N$ and the length of the original sequence $L$. The length of the remaining space is equal to $N-L-1$ if delimiter 0 is used, otherwise $N-L$. Let’s take the case of using the delimiter 0 as an example. Next, we randomly truncate the sub-sequence from the original sequence based on this length:
(6) $s_{u}^{sub}=subseq\left(s_{u},N-L-1\right).$
Then, we use $s_{u}^{sub}$ as the padding content:
[…]
It should be noted that the above operation is only applicable to sequences where the remaining space is insufficient for padding the whole original sequence, and the delimiter 0 is optimal. As illustrated in Figure 2 (f), this padding operation is only applied to sequence $[0,0,0,0,v_{6},v_{4},v_{1},v_{7},v_{9},v_{2}]$. Applying Eq. 6 we get sub-sequence $[v_{4},v_{1},v_{7}]$. Through Eq. 7, the final padded sequence is $[v_{4},v_{1},v_{7},0,v_{6},v_{4},v_{1},v_{7},v_{9},v_{2}]$ through Eq. 7.
Based on the RepPad in Section 3.4, we add this operation and refer to this version as RepPad+. As mentioned in Section 3.4, in practice, we strongly recommend users apply function Eq. (2) after using our method, which can be seen as a final check on sequence length. Due to the influence of different training data and code implementations, the length of the augmented sequence may not be $N$, and using Eq. (2) can ensure that the length of the sequence meets the conditions to be fed into the model. We present the Python procedure of RepPad+ in Algorithm 2. Note that for RepPad+, since medium-length sequences are not suitable for full repeated padding, the number of $m$ has no effect on their augmentation. Only whether or not the delimiter 0 is added (i.e., whether there is a delimiter 0 between the padding sub-sequence and the original sequence) will have an effect on the augmentation of medium-length sequences.
Algorithm 2 RepPad+ Pseudocode ($m=random(1,max)$)
# orig_seq: An Original User Sequence s_u.
# max_len: Maximum Sequence Length N.
def reppad+(orig_seq, max_len):
final_input = orig_seq[:-3] # training items
final_target = orig_seq[1:-2] # target items
if len(orig_seq[:-3]) > max_len - 2:
# No space for padding operation. One position for delimiter 0, One position for an item.
return final_input, final_target
elif int(max_len / len(final_input)) <= 1:
# Not enough space left for padding whole original sequence. Padding sub-sequence.
sub_sequence_len = max_len - 1 - len(final_input)
start_index = random.randint(0, len(final_input) - sub_sequence_len)
final_input = final_input[start_index:start_index + sub_sequence_len] + [0] + final_input
final_target = final_target[start_index:start_index + sub_sequence_len] + [0] + final_target
return final_input, final_target
else:
# Padding original sequence.
# Calculate the maximum padding count.
max_pad_num = int(max_len / len(final_input))
# Randomly select a padding count.
pad_num = random.randint(1, max_pad_num)
# Perform the Repeated Padding operation.
final_input = (final_input + [0]) * pad_num + final_input
final_target = (final_target + [0]) * pad_num + final_target
return final_input, final_target
# Then apply function Eq.(2) before being fed into the model.
3.6. Discussions
3.6.1. Comparison with Existing Methods.
We analyze the advantages of our approach over existing methods from the following perspectives:
  • (i) No increase in training data size: From Figure 2, we can observe that our RepPad+ performs data augmentation for both short and medium sequences without increasing the amount of data for training. While previous approaches can be viewed as extending the length of the dataset vertically, our approach utilizes the idle space horizontally. For traditional data augmentation methods (Tan et al., 2016; Tang and Wang, 2018; Wang et al., 2021), each sequence will be augmented at least once in each epoch, so the amount of data will be at least $2\left|\mathcal{U}\right|$. Some methods augment specific sequences (e.g., cold start user sequence) in the dataset (Liu et al., 2021b; Wang et al., 2024, 2022), so the total number of sequences will be at least greater than $\mathcal{U}$. For sequential recommendation models that leverage both data augmentation and contrastive learning (Xie et al., 2022; Liu et al., 2021a; Dang et al., 2023b), each sequence is augmented at least twice, and with the addition of the original sequence, the number of model sequences involved in training is at least $3\left|\mathcal{U}\right|$. In addition, according to the optimal parameter settings of some methods (Liu et al., 2023; Yin et al., 2024), the total number of sequences in a single epoch may reach $4\left|\mathcal{U}\right|$ or $5\left|\mathcal{U}\right|$.
  • (ii) No hyper-parameters and no increase in model parameters: These existing augmentation methods or augmentation operators (Slide Window, Dropout, Crop, Insert, Mask, Substitute, Reorder, and so on) typically contain at least one hyper-parameter for each that must be manually tuned based on the dataset (Tan et al., 2016; Tang and Wang, 2018; Xie et al., 2022; Liu et al., 2021a). Some training-required data generation modules introduce additional model parameters in addition to hyper-parameters (Wang et al., 2022; Liu et al., 2023; Yin et al., 2024). However, our approach does not involve any hyper-parameters that need to be tuned carefully, nor does it increase the number of parameters in the original model.
  • (iii) No training cost and model agnostic: Some researchers propose to use bidirectional Transformers (Liu et al., 2021b; Jiang et al., 2021), diffusion models (Liu et al., 2023) or encoder-decoder based augmenter (Yin et al., 2024) for data augmentation. These methods require training of data augmentation modules in addition to training of backbone recommendation networks and have limitations on the type of backbone network. In contrast, our method does not require extra training processes and can be adapted to most sequential models. We present a summary comparison between PepPad+ and existing methods in Table 2. It should be emphasized that our focus is on sequence-level data augmentation, which is also the most mainstream practice in data augmentation in the field of sequence recommendation. Some works explore augmentation at the embedding (Qiu et al., 2022) and model (Hao et al., 2023) levels. These methods do not operate directly on the sequence data and are not involved in our discussion.
Table 2. Comparison of sequence-level data augmentation between our method and several representative methods. The ‘No HP & MP’ represents no additional hyper-parameters or increases in model parameters. The ‘No Training’ represents no training process for data generation modules or auxiliary tasks. The ‘TTS’ represents the total training sequences in each epoch.
Methods Augmentation Type No HP & MP No Training Model Agnostic TTS
Only Augmentation (Tan et al., 2016; Tang and Wang, 2018) Random or Elaborate Operators $\geq$ $2\left|\mathcal{U}\right|$
ASReP (Liu et al., 2021b) and BiCAT (Jiang et al., 2021) Bidirectional Transformers $>$ $\left|\mathcal{U}\right|$
LLM Augmentation (Wang et al., 2024) Large Language Models $>$ $\left|\mathcal{U}\right|$
RepPad (Dang et al., 2024) Repeated Padding for short $=$ $\left|\mathcal{U}\right|$
RepPad+ (Ours) Repeated Padding for short and medium $=$ $\left|\mathcal{U}\right|$
@@ 3.6. Discussions
Our approach intends to maximize the use of idle input space. Intuitively, it is more effective on short-sequence datasets and may be ineffective on long-sequence datasets. In long-sequence datasets, most users have sequence lengths that close to or exceed the given maximum sequence length, and thus, there will be no space left for repeated padding. For RepPad and our proposed RepPad+, the latter will outperform the former on long sequence datasets. This is because RepPad+ performs augmentation on both short and medium-length sequences, whereas the RepPad performs only on short sequences. As shown in Figure 1 (a), only a very small fraction of users (0.72% in the Beauty dataset) have sequences longer than the maximum sequence length $N$. We will verify our conjecture in the experimental section.
@@ 4.1. Experiment Setup
4.1.1. Datasets
The experiments are conducted on seven widely-used public datasets:
@@ 4.1.1. Datasets
  • Toys, Beauty, Sports, and Home: Four datasets obtained from Amazon (McAuley et al., 2015) and correspond to the ”Toys and Games”, ”All Beauty, ”Sports and Outdoors”, and ”Home and Kitchen” categories, respectively.
  • Yelp [^1] https://www.yelp.com/dataset: This is a popular dataset for business recommendation. As it is very large, we only use the transaction records after January 1st, 2019.
  • LastFM [^2] https://grouplens.org/datasets/hetrec-2011/: It is a music artist recommendation dataset and contains user tagging behaviors for artists.
@@
  • MovieLens-1M [^3] https://grouplens.org/datasets/movielens/ (ML-1M for short): It comprises rating behaviors gathered from a movie review website.
@@
The Amazon and Yelp datasets contain mainly short user sequences. The MovieLens-1M and LastFM datasets contain mainly long user sequences. Following (Liu et al., 2021a; Dang et al., 2023a; Tian et al., 2023), we reduce the data by extracting the 5-core. The maximum sequence length $N$ is set to 200 for the MovieLens-1M dataset, 100 for LastFM dataset, and 50 for other datasets. The detailed statistics are summarized in Table 3.
Table 3. The statistics of seven datasets. The ‘AL’ denotes average length.
Dataset # Users # Items # Interactions # AL Sparsity
LastFm 1,090 3,646 52,551 48.2 98.68%
Toys 19,412 11,924 167,597 8.6 99.93%
Beauty 22,363 12,101 198,502 8.9 99.93%
Sports 35,598 18,357 296,337 8.3 99.95%
Yelp 30,431 20,033 316,354 10.4 99.95%
Home 66,519 28,237 551,682 8.3 99.97%
ML-1M 6,040 3,706 1,000,209 165.6 95.53%
4.1.2. Baselines
@@
To provide a comprehensive evaluation of RepPad+, the baselines consist of three main categories:
@@
First category is eight representative sequential recommendation models, which are used to evaluate whether RepPad+ is effective in improving the model’s performance.
@@
  • RNN-based models: GRU4Rec (Hidasi et al., 2015) leverages gated recurrent units to model behavioral patterns. It argues that by modeling the whole session, more accurate recommendations can be provided. NARM (Li et al., 2017) incorporates attention mechanism into RNN architectures. It proposes to capture both the user’s sequential behavior and main purpose in the current session.
@@
  • CNN-based models: Caser (Tang and Wang, 2018) models recent actions as an ”image” among time and latent dimensions. It learns sequential patterns using convolutional filters. NextItNet (Yuan et al., 2019) combines masked filters with 1D dilated convolutions to increase the receptive fields, which is important to model the long-range dependencies.
  • MLP-based models: gMLP suggests a simpler alternative to the multi-head self-attention layers in Transformers. It adopts a cross-token MLP structure named Spatial Gating Unit to capture sequential information. FMLP-Rec (Zhou et al., 2022) finds that the logged user behavior data usually contains noisy interactions, and filtering algorithms from the digital signal processing area are useful to alleviate the influence of the noise in deep SR models. Inspired by this finding, an all-MLP model with learnable filters for SR tasks is proposed.
  • Transformer-based models: SASRec (Kang and McAuley, 2018) is the representative work in a sequential recommendation that leverages the multi-head attention mechanism to the next item recommendation. It models the entire user sequence (without any recurrent or convolutional operations). LightSANs (Fan et al., 2021) introduces the low-rank decomposed self-attention, which projects the user’s historical items into a small constant number of latent interests and leverages item to-interest interaction to generate the context-aware representation. They propose overcoming quadratic complexity and over-parameterization.
@@
Second category is the widely used stochastic data augmentation methods, which does not require training but include several hyper-parameters that need to be manually tuned.
@@
  • Slide Window (SW) (Tang and Wang, 2018): It adopts slide windows to intercept multiple new subsequences from the original data.
@@
  • CMR (Xie et al., 2022): It proposes three augmentation operators with contrastive learning. In this category, we only use the three operators, including Crop, Mask, and Reorder.
@@
  • CMRSI (Liu et al., 2021a): Based on CMR, this work proposes two informative operators. We adopt five operators including Crop, Mask, Reorder, Substitute, and Insert.
@@
Third category is representative of training-required augmentation models, which usually include trainable parameters and hyper-parameters that need to be manually tuned.
@@
  • ASReP (Liu et al., 2021b): It employs a reversely pre-trained transformer to generate pseudo-prior items for short interaction sequences. Then, fine-tune the pre-trained transformer to predict the next item.
@@
  • DiffuASR (Liu et al., 2023): It adopts the diffusion model for discrete sequence generation tasks. Besides, two guide strategies are designed to make the generated items correspond to the raw data.
@@
  • CL4SRec (Xie et al., 2022): This method leverages random data augmentation operators with contrastive learning to extract self-supervised signals from the original data.
@@
We do not include methods L2Aug (Wang et al., 2022), CASR (Wang et al., 2021), BiCAT (Jiang et al., 2021), and LLM Augmentation (Wang et al., 2024) since they do not provide open-source codes for reliable reproduction.
4.1.3. Evaluation Metrics
@@
Following (Liu et al., 2021a; Dang et al., 2023a; Tian et al., 2023), we adopt the leave-one-out strategy, wherein the last item of each user sequence serves as the test data, the items preceding it as validation data, and the remaining data as training data. We rank the prediction over the whole item set rather than negative sampling, otherwise leading to biased discoveries (Krichene and Rendle, 2020). The evaluation metrics include Hit Ratio@K (denoted by H@K), and Normalized Discounted Cumulative Gain@K (N@K). We report results with K $\in\{5,10,20\}$. Generally, greater values imply better ranking accuracy.
4.1.4. Implementation Details